One company wants to start a new project At first, company put 5.000.000 EUR The net working capital requirements are equal to 15% of the annual sales at the beginning of each year The selling price of the product is 20,000 per unit for first year and the variable costs for the first year are 15,000 per unit and fixed costs for first year 1.000.000 EUR million After first year, price, variable and fixed costs will increase with inflation rate We also know that a)inflation rate is 2% and projects life is 4 years. b) Depreciations rates per year 15%,18%,11% and 14%. After 4 years the market value of the equipment will cost 420.000€ and the cost of capital is 10%

Tax rate is 20% and the estimated sales per year 1000 units.

How can we compute NPP,IRR and payback?
Answer #1
0 1 2 3 4
Sales [1000 units*$20000+inflation of 2% from 2nd year] 20000000 20400000 20808000 21224160
Variable cost [15000 per unit + inflation of 2% from 2nd year] 15000000 15300000 15606000 15918120
Fixed costs [1000000 in the first year + inflation of 2%] 1000000 1020000 1040400 1061208
Depreciation 750000 900000 550000 700000 2900000
EBIT 3250000 3180000 3611600 3544832
Tax at 20% 650000 636000 722320 708966
NOPAT 2600000 2544000 2889280 2835866
Add: Depreciation 750000 900000 550000 700000
OCF 3350000 3444000 3439280 3535866
Capital expenditure 5000000
After tax salvage value [see note at the end] 756000
Change in NWC 3000000 60000 61200 62424 -3183624
Annual after tax project cash flows -8000000 3290000 3382800 3376856 7475490
After tax salvage value:
Salvage value 420000
Book value = 5000000-2900000 = 2100000
Loss on sale = 2100000-420000 = 1680000
Tax shield on loss at 20% 336000
After tax salvage value 756000
NPV:
After tax project cash flows -8000000 3290000 3382800 3376856 7475490
PVIF at 10% 1 0.90909 0.82645 0.75131 0.68301
PV at 10% -8000000 2990909 2795702 2537082 5105860
NPV 5429553
IRR:
IRR is that discount rate for which NPV = 0. This has to be found out by trial and error to get 0 NPV.
Discounting with 30% NPV
PVIF at 35% 1 0.74074 0.54870 0.40644 0.30107
PV at 35% -8000000 2437037 1856132 1372496 2250632 -83702
PVIF at 34% 1 0.74627 0.55692 0.41561 0.31016
PV at 34% -8000000 2455224 1883939 1403454 2318571 61188
IRR = 34+61188/(83702+61188) = 34.42%
PAYBACK:
Cumulative Cash Flows -8000000 -4710000 -1327200 2049656 9525146
Payback period = 2+1327200/3376856 = 2.39 Years